OpenStudy (anonymous):

how do you simplify: 12.5t-2.5-3t+4-3.5t

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

combine like terms first!

OpenStudy (anonymous):

ok, so the like terms are: 12.5t, 3t, and 3.5t

OpenStudy (anonymous):

so 9t

OpenStudy (anonymous):

yes, so you would get? 12.5t+3t+3.5t=6t! then you combine the other like terms so -2.5+4=?

OpenStudy (anonymous):

oops, I meant 19t

OpenStudy (anonymous):

I mean 12.5t-3t-3.5t=6t how'd you get 19t?

OpenStudy (anonymous):

I added the like terms together

OpenStudy (anonymous):

there's a - in front of 3.5t and 3t so you subtract those from 12.5t got it?

OpenStudy (anonymous):

yah ok, so its 6t

OpenStudy (anonymous):

so then you would simplifies the others so -2.5+4=?

OpenStudy (anonymous):

6.5

OpenStudy (anonymous):

nope remember 2.5 is negative so it basically the same as 4-2.5=?

OpenStudy (anonymous):

oops forgot to put in the negative, its really 1.5

OpenStudy (anonymous):

yes!!! so your new equation would be 6t+1.5!

OpenStudy (anonymous):

so, 7.5t

OpenStudy (anonymous):

No that is your simplified answer lol 6t+1.5 because there are no more like terms!! hope this helps:)

OpenStudy (anonymous):

ok, so 6t+1.5 is the answer

OpenStudy (anonymous):

Your like terms always come first no matter what then you have to get to letters by themselves and after you combine like terms you add them to the rest of the equation

OpenStudy (anonymous):

thanks